OpenStudy (anonymous):

Use the right triangle to solve the triangle. A=40 degrees, B=50 degrees, c=10 Find b,a

OpenStudy (anonymous):



Nnesha (nnesha):

 use \[\rm sin = \frac{ opposite }{ hypotenuse }~~~~ \cos = \frac{ adjacent }{ hypotenuse } ~~\tan= \frac{ opposite }{ adjacent }\]